Find a poner series representation for the function and determine the interval of convergence f(x) It 4x²
Answer #1

Solution: 2x a) Let 1(x) = 1+4x* = 2x (14) 3 1+4x? 1 1 Consider the function : 1+47 We have power series for -=1+x+x2 + 1-X S